2. Saving and investment in the national income accounts The following table contains data for a hypothetical closed economy that uses the dollar as its currency. Suppose GDP in this country is $1,680 million. Enter the amount for government purchases. Value (Millions of dollars,) National Income Account Government Purchases (G) Taxes minus Transfer Payments (T) Consumption (C) Investment (I) 360 1,000 280

The options are : National savings = (Y - C - G) or (Y-C) or (G-T) or (Y-T-G)

for the second blank under National Savings the options are (Y) or (I) or (C) or (G)

The options are : Private savings = (Y - C - T) or (Y - T - I) or (T - G) or (C -T)

The options are : Public Savungs = (Y - C - T) or (Y - T - I) or (T - G) or (C -T)

Based on your calculations, the government is running a budget (Surplus , Deficit)
Answer #1

Calculate Government purchases (G) -

G = GDP - Consumption - Investment

G = $1,680 million - $1,000 million - $280 million

G = $400 million

The government purchases (G) is $400 million.

Calculate National Saving -

National Saving (S) = Y - C - G

National Saving (S) = I

National Saving (S) = $280 million

Calculate Private Saving -

Private Saving = Y - C - T

Private Saving = $1,680 million - $1,000 million - $360 million

Private Saving = $320 million

Calculate the Public Saving -

Public Saving = T - G

Public Saving = $360 million - $400 million

Public Saving = -$40 million

The negative value of public saving implies deficit.

So,

The government is running a budget deficit.